OpenStudy (anonymous):

have a question........ Input the equation of the given line in standard form. The line with m =-1/2 and the point (-2, -3).

OpenStudy (anonymous):

y-(-3)= -1/2 (x-(-2))

OpenStudy (anonymous):

y+3=1/2x+1????????

OpenStudy (anonymous):

-{(x+2)}/2

OpenStudy (***[isuru]***):

@divu.mkr u have write it in the point slope form it;s not the slandered form

OpenStudy (anonymous):

ok... if you want point slope form just subtract both sides by 3

OpenStudy (***[isuru]***):

the slandered form is represent by the format y = mx + c where m = slope c = y - intercept now the question has given the slope.... so we can write it as y = -(1/2)x + c now the other data given is a coordinate of a point (-2 , -3 ) which mean in the given point x = -2 y = -3 so .. u can plug them in to the function y = -(1/2)x + c -3 = -(1/2)(-2) + c - 3 = +1 + c c = -4 so.. the function of the line in the slandered format will be.. y = -(1/2)x -4 got it :-) ?

OpenStudy (anonymous):

ya...thankx

OpenStudy (***[isuru]***):

u r welcome!